OpenStudy (anonymous):

vThree charged particles lie on the x-axis. A particle with a charge q1 = 13.62 x 10-6 C is at x = 2.60 m and a particle with a charge of q2 = 6.80 x 10-6 C is at the origin. Where on the x-axis can a negative charge, q3, be placed such that the resultant force on it is zero?

OpenStudy (anonymous):

I know coulomb's law, tried adding the force from Q1Q3 and Q2Q3 to equal zero

OpenStudy (anonymous):

The force is zero where the electric field is zero, and that's the location where (if a=2.6 m) \[ \frac{Q1}{(x-a)^2} = \frac{Q2}{x^2} \]